Suppose that the speed v (in ft/s) of a skydiver t seconds after leaping from a plane is given by the equation ๐‘ฃ = 190(1 โˆ’ ๐‘’ โˆ’ 0.168๐‘ก).

(a) Graph ๐‘ฃ versus ๐‘ก.

(b) By evaluating an appropriate limit, show that the graph of ๐‘ฃ versus ๐‘ก has a horizontal asymptote ๐‘ฃ = ๐‘ for an appropriate constant ๐‘.

(c) What is the physical significance of the constant ๐‘ in part (b)?


1
Expert's answer
2021-10-04T11:40:50-0400

(a) Graph ๐‘ฃ versus ๐‘ก.





Once we graph this we find that there is an asymptoteย for v = c = 190.


(b) By evaluating an appropriate limit, show that the graph of ๐‘ฃ versus ๐‘ก has a horizontal asymptote ๐‘ฃ = ๐‘ for an appropriate constant ๐‘.


We have to find the limit when "t\to \infin" so


"\lim\limits_{x\, \to \infin} v=\lim\limits_{x\, \to \infin} 190(1-e^{-0.168t})
\\ \lim\limits_{x\, \to \infin} 190(1-e^{-0.168t})=190-\lim\limits_{x\, \to \infin} e^{-0.168t}
\\ \lim\limits_{x\, \to \infin} 190(1-e^{-0.168t})=190"


(c) What is the physical significance of the constant ๐‘ in part (b)?


c = 190 is the velocity once "t\to \infin", and this will be the limit speed expressed in ft/s.



Reference:

  • Thomas, G. B., & Finney, R. L. (1961).ย Calculus. Addison-Wesley Publishing Company.